OpenStudy (anonymous):

Absolute value equation help! |x-1|=5x+10

OpenStudy (jdoe0001):

\(\large |x-1|=5x+10\implies \begin{cases} +(x-1)=5x+10\implies &x=?\\ -(x-1)=5x+10\implies &x=? \end{cases}\)

OpenStudy (zzr0ck3r):

\[|x-1|=5x+10\] means two things 1) \(x-1=5x+10\) 2) \(-(x-1)=5x+10\) Solve both of these to get two answers.